Chem help Partial Pressure?
I was hoping someone can help me figure out how this Partial Pressure/Mole fraction problem is solved. I know the answer is 4.2 atm but Im not quite sure how that answer is derived.
A diver breathes a heliox mixture with an oxygen mole fraction of 0.050. What must the total pressure be for the partial pressure of oxygen to be 0.21 atm?
Thank you!

Grade:12

2 Answers

Mujahid Ahmed
10 years ago

Partial pressure of gas = Mole fraction x Total pressure.

0.21 = 1/20 x P

P = 4.2 atm

 

Approve if i helped :)

Vikas TU
10 years ago

From Gauss''s law it is:
partial pressure of oxygen = mole fraction of O2 x total pressure
or
Pa = Xa.Pt

0.21 = 0.050 x Pt
Pt = 0.21/0.050
= 4.2 atm

Derivation:

Gauss''s laws states that in gaseous medium or .ie. if the mixture having all in gaseous state then some conditions have been brought under practical observation. And u will need to remind them for doing further calculations.

1) Partial pressure of each gas or component is given as:

Pa = Xa . Pt (say component a)
or
Pb = Xb .Pt (say component b)

2) Total pressure is equals to sum of all the partial pressures of all components (gases)

Pt = Pa + Pb (always remember)

If u sort out this:

U will also get an relation as:

Xa + Xb = 1
which says that sum of the mol fractions of all component is always equals to 1.

 

Plz  Approve!